# theoremTheorem

### v2

@@ Abstract.
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXconference: ACM FAccT 2025; June 23–26, 2025; Athens, Greececopyright: none
@@ 2. Related Work
While this approach could achieve lower prices for consumers if deployed, this assumes the ability and willingness of a seller to do so. To mitigate this issue, we design a solution to unfair pricing without directly requiring cooperation from the seller. Some work has taken this approach—counterfactual fairness (Kusner2017) could allow an individual to achieve better pricing without direct access to the underlying pricing system. However, this requires both an accurate causal model and toggles that individuals can act on (10.1145/3442188.3445899). On the collective action side, hardt2023algorithmic quantified the effect of coordinated feature changes against an algorithmic system. In our work, collaboration does not aim to change the direct outcome of an algorithmic system, but rather to adjust final outcomes for individuals. Our work does not require a causal model or online behavior change for individuals to benefit. Rather, we focus on consumers who are incentivized by monetary reward to share information. We specifically focus on cases where the goods are identical and purchased on online marketplaces, where potentially unfair differential pricing can occur without users’ knowledge (Lippert2023; miller2014we).
@@ 3. Problem Statement
We frame the matching process $\mathcal{P}$ as a network problem. Our consumers exist in a directed network $G=(V,E)$ where $(u,v)\in E$ if $p_{u}>p_{v}$. Interactions $\mathcal{J}$ are matched directed edges; each consumer $v$ has at most $k$ matched incoming edges. The outgoing edge count is capped at 1, since each consumer can only buy the good $g$ once. This resource constraint $k$ represents how much time a consumer is willing to spend on this system. If an interaction $j_{uv}$ is executed, the edge $(u,v)$ has been transacted on. We call $u$ the buyer and $v$ intermediary; one consumer might be matched as both intermediary and buyer on different transactions.
In this paper we examine which design of exchange system $\mathcal{S}$ will maximally improve our fairness targets $\mathcal{F}$ on market $\mathcal{M}$, and under what circumstances it is financially feasible to maintain. Specifically we ask:
@@ 6.2. RQ 2: Feasibility of our fairness measures
In Figure 2 we show how optimizing for each fairness target $\mathcal{F}$ affects the measure of each outcome. Here we will focus on the decentralized scenario, as centralized price-setting results in minimal trades. We see that while specifically optimizing for $\mu_{I}$ and $\mu_{\mathcal{G}}$, decentralized price-setting is able to effectively reduce the average price paid. On the other hand, achieving low standard deviation (Figures 2(b) and 2(d)) is more difficult. We can investigate how optimizing for each fairness metric compares to the “ideal” scenario. In Table 2 we show numerically that $\mu$-optimizing procedures are more feasible under decentralized settings, while $\sigma$-optimizing are less so. During decentralized negotiations, as $k$ (the number of times a consumer can act as intermediary) increases, intermediaries are able to benefit multiple times while buyers only benefit once. This increases the gap, even when choosing $\sigma$-optimizing pairings.
@@ 7. Discussion
If we allow for a repeated game, there are many ways a marketplace might respond: 1) increasing technical barriers for the consumer that block trades, 2) increasing personalization, 3) decreasing personalization but increasing prices overall. While we cannot predict which of these might occur, the modeling of our system is robust to all of these changes–increased technical barriers can be modeled as additional disutility to each consumer from participating in each transaction; increased personalization would actually increase the fairness of outcomes; decreased personalization, while decreasing fairness, can be easily modeled with a change to pricing algorithm $\mathcal{A}$.
Another issue that may arise between the system and market is that owners of the exchange system could forgo fairness and accountability in exchange for higher revenue. If the system and market collude, the market could produce prices that when passed through the colluding exchange system $\mathcal{S}$, result in worse outcomes for consumers, but higher revenue for the system and market. Considering our finding that higher dispersion can increase revenue and decrease net cost to consumers, existence of such a strategy is possible. Thus, the assumption that the system is independent from the market and has different goals is key.
@@ 8. Future Work and Limitations
Human studies: In this work we make simplifying assumptions regarding consumers’ actions, e.g., the form of utility functions, willingness to participate, demographic attributes, and resource constraints. To test these assumptions, future human studies would be ideal. This would allow us to test our design principles on users who may not respond as we originally modeled them. Other design principles could be tested as well (e.g., wording of messaging to consumers) to incentivize trading.

### v3

@@ 2. Related Work
While this approach could achieve lower prices for consumers if deployed, this assumes the ability and willingness of a seller to do so. To mitigate this issue, we design a solution to unfair pricing without directly requiring cooperation from the seller. Some work has taken this approach—counterfactual fairness (Kusner2017) could allow an individual to achieve better pricing without direct access to the underlying pricing system. However, this requires both an accurate causal model and toggles that individuals can act on (10.1145/3442188.3445899). On the collective action side, hardt2023algorithmic quantified the effect of coordinated feature changes against an algorithmic system. In our work, collaboration does not aim to change the direct outcome of an algorithmic system, but rather to adjust final outcomes for individuals. Our work does not require a causal model or online behavior change for individuals to benefit. Rather, we focus on consumers who are incentivized by monetary reward to share information. We specifically are interested in cases where the goods are identical and purchased on online marketplaces, where potentially unfair differential pricing can occur without users’ knowledge (Lippert2023; miller2014we).
@@ 3. Problem Statement
We frame the matching process $\mathcal{P}$ as a network problem. Our consumers exist in a directed network $G=(V,E)$ where $(u,v)\in E$ if $p_{u}>p_{v}$. Interactions $\mathcal{J}$ are matched directed edges; each consumer $v$ has at most $k$ matched incoming edges. The outgoing edge count is capped at 1, since each consumer can only buy the good $g$ once. This resource constraint $k$ represents how much time a consumer is willing to spend on this system. If an interaction $j_{uv}$ is executed, the edge $(u,v)$ has been transacted on. We call $u$ the buyer and $v$ intermediary; one consumer might be matched as both an intermediary and a buyer on different transactions.
In this paper, we examine which design of exchange system $\mathcal{S}$ will maximally improve our fairness targets $\mathcal{F}$ on market $\mathcal{M}$, and under what circumstances it is financially feasible to maintain. Specifically we ask:
@@ 6.2. RQ 2: Feasibility of our fairness measures
In Figure 2 we show how optimizing for each fairness target $\mathcal{F}$ affects the measure of each outcome. Here we will focus on the decentralized scenario, as centralized price-setting results in minimal trades. We see that when specifically optimizing for $\mu_{I}$ and $\mu_{\mathcal{G}}$, decentralized price-setting is able to effectively reduce the average price paid. On the other hand, achieving low standard deviation (Figures 2(b) and 2(d)) is more difficult. We can investigate how optimizing for each fairness metric compares to the “ideal” scenario. In Table 2 we show numerically that $\mu$-optimizing procedures are more feasible under decentralized settings, while $\sigma$-optimizing are less so. During decentralized negotiations, as $k$ (the number of times a consumer can act as intermediary) increases, intermediaries are able to benefit multiple times while buyers only benefit once. This increases the gap, even when choosing $\sigma$-optimizing pairings.
@@ 7. Discussion
If we allow for a repeated game, there are several ways a marketplace might respond: 1) increasing technical barriers for the consumer that block trades, 2) increasing personalization, 3) decreasing personalization but increasing prices overall. While we cannot predict which of these might occur, the modeling of our system is robust to all of these changes–increased technical barriers can be modeled as additional disutility to each consumer from participating in each transaction; increased personalization would actually increase the fairness of outcomes; decreased personalization, while decreasing fairness, can be easily modeled with a change to pricing algorithm $\mathcal{A}$.
Another issue that may arise between the system and market is that owners of the exchange system could forgo fairness and accountability in exchange for higher revenue. If the system and market collude, the market could produce prices that, when passed through the colluding exchange system $\mathcal{S}$, result in worse outcomes for consumers, but higher revenue for the system and market. Considering our finding that higher dispersion can increase revenue and decrease net cost to consumers, existence of such a strategy is possible. Thus, the assumption that the system is independent from the market and has different goals is key.
@@ 8. Future Work and Limitations
Human studies: In this work we make simplifying assumptions regarding consumers’ actions, e.g., the form of their utility functions, willingness to participate, demographic attributes, and resource constraints. To test these assumptions, future human studies would be ideal. This would allow us to test our design principles on users who may not respond as we originally modeled them. Other design principles could be tested as well (e.g., wording of messaging to consumers) to incentivize trading.